# Thermodynamic nonequilibrium effects in three-dimensional high-speed compressible flows: Multiscale modeling and simulation via the discrete Boltzmann method

### v1

@@ Acknowledgements.
The authors sincerely thank Zhaowen Zhuang, Jiahui Song for helpful discussions. We acknowledge support from the National Natural Science Foundation of China (Grant Nos. U2242214,11875001 and 12172061), Hebei Outstanding Youth Science Foundation (Grant No. A2023409003), Central Guidance on Local Science and Technology Development Fund of Hebei Province (Grant No. 226Z7601G), Fujian Provincial Units Special Funds for Education and Research (Grant No. 2022639), the Foundation of National Key Laboratory of Shock Wave and Detonation Physics (Grant no. JCKYS2023212003).
@@ Appendix A Second-order viscous stress and heat flux in three dimensions
The expressions for the second-order viscous stress $\bm{\Delta}_{2}^{*(2)}$ and heat flux $\bm{\Delta}_{31}^{*(2)}$ in three-dimensions are presented in Table 1, with the associated terms $S_{1},\dots,S_{36}$ listed in Table LABEL:TableII.
Table 2: Expressions for $\bm{\Delta}_{2}^{*(2)}$ and $\bm{\Delta}_{31}^{*(2)}$.
@@
$\Delta_{2xx}^{*(2)}$ $-2n_{3}^{-2}\tau^{2}(R^{2}T^{2}n_{3}S_{1}-\rho RTn_{-1}S_{2}+\rho RTS_{3}-R^{2}T^{2}\frac{S_{4}}{\rho}-\rho R^{2}S_{5})$ non-organized momentum flux (NOMF)/generalized viscous stress
@@
(3) Enhanced cross-coupling among driving forces: The cross-coupling among nonequilibrium driving forces becomes highly intricate, involving terms such as: $\partial_{x}u_{x}\partial_{z}u_{z}$, $\partial_{y}u_{x}\partial_{z}u_{z}$, $\partial_{z}u_{x}\partial_{z}u_{z}$, $\partial_{x}u_{y}\partial_{z}u_{z}$, $\partial_{y}u_{y}\partial_{z}u_{z}$, $\partial_{z}u_{y}\partial_{z}u_{z}$, $\partial_{x}u_{z}\partial_{z}u_{z}$, $\partial_{y}u_{z}\partial_{z}u_{z}$, $\partial_{x}u_{x}\partial_{z}u_{x}$, $\partial_{y}u_{y}\partial_{z}u_{x}$, $\partial_{z}u_{y}\partial_{z}u_{x}$, $\partial_{x}u_{x}\partial_{z}u_{y}$, $\partial_{z}u_{x}\partial_{z}u_{y}$, $\partial_{y}u_{y}\partial_{z}u_{y}$, $\partial_{x}u_{x}\partial_{x}u_{z}$, $\partial_{x}u_{y}\partial_{x}u_{z}$, $\partial_{y}u_{y}\partial_{x}u_{z}$, $\partial_{y}u_{x}\partial_{y}u_{z}$, $\partial_{x}\rho\partial_{z}\rho$, $\partial_{y}\rho\partial_{z}\rho$, $\partial_{x}T\partial_{z}T$, $\partial_{y}T\partial_{z}T$, $\partial_{z}u_{x}\partial_{x}T$, $\partial_{x}u_{z}\partial_{x}T$, $\partial_{z}u_{y}\partial_{y}T$, $\partial_{y}u_{z}\partial_{y}T$, $\partial_{z}u_{z}\partial_{y}T$, $\partial_{x}u_{x}\partial_{z}T$, $\partial_{z}u_{x}\partial_{z}T$, $\partial_{y}u_{y}\partial_{z}T$, $\partial_{z}u_{y}\partial_{z}T$, $\partial_{x}u_{z}\partial_{z}T$, $\partial_{y}u_{z}\partial_{z}T$.

### v2

@@ Acknowledgements.
The authors sincerely thank Zhaowen Zhuang, Jiahui Song for helpful discussions. We acknowledge support from the National Natural Science Foundation of China (Grant Nos. U2242214, 52278119, 11875001 and 12172061), Hebei Outstanding Youth Science Foundation (Grant No. A2023409003), Central Guidance on Local Science and Technology Development Fund of Hebei Province (Grant No. 226Z7601G), Fujian Provincial Units Special Funds for Education and Research (Grant No. 2022639), the Foundation of National Key Laboratory of Shock Wave and Detonation Physics (Grant no. JCKYS2023212003), the Opening Project of State Key Laboratory of Explosion Science and Safety Protection (Beijing Institute of Technology) (Grant No. KFJJ25-02M), and science foundation of NCIAE (Grant No. ZD-2025-06).
@@ Appendix A Second-order viscous stress and heat flux in three dimensions
The expressions for the second-order viscous stress $\bm{\Delta}_{2}^{*(2)}$ and heat flux $\bm{\Delta}_{3,1}^{*(2)}$ in three-dimensions are presented in Table LABEL:TableII, with the associated terms $S_{1},\dots,S_{36}$ listed in Table LABEL:TableIII.
Table 2: Expressions for $\bm{\Delta}_{2}^{*(2)}$ and $\bm{\Delta}_{3,1}^{*(2)}$.
@@
$\Delta_{2xx}^{*(2)}$ $-2n_{3}^{-2}\tau^{2}(n_{3}R^{2}T^{2}S_{1}-n_{-1}\rho RTS_{2}+\rho RTS_{3}-R^{2}T^{2}\frac{S_{4}}{\rho}-\rho R^{2}S_{5})$ non-organized momentum flux (NOMF)/generalized viscous stress
@@
(3) Enhanced cross-coupling among driving forces: The cross-coupling among nonequilibrium driving forces becomes highly intricate, involving terms such as: $\partial_{x}u_{x}\partial_{z}u_{z}$, $\partial_{y}u_{x}\partial_{z}u_{z}$, $\partial_{z}u_{x}\partial_{z}u_{z}$, $\partial_{x}u_{y}\partial_{z}u_{z}$, $\partial_{y}u_{y}\partial_{z}u_{z}$, $\partial_{z}u_{y}\partial_{z}u_{z}$, $\partial_{x}u_{z}\partial_{z}u_{z}$, $\partial_{y}u_{z}\partial_{z}u_{z}$, $\partial_{x}u_{x}\partial_{z}u_{x}$, $\partial_{y}u_{y}\partial_{z}u_{x}$, $\partial_{z}u_{y}\partial_{z}u_{x}$, $\partial_{x}u_{x}\partial_{z}u_{y}$, $\partial_{y}u_{y}\partial_{z}u_{y}$, $\partial_{x}u_{x}\partial_{x}u_{z}$, $\partial_{x}u_{y}\partial_{x}u_{z}$, $\partial_{y}u_{y}\partial_{x}u_{z}$, $\partial_{y}u_{x}\partial_{y}u_{z}$, $\partial_{x}\rho\partial_{z}\rho$, $\partial_{y}\rho\partial_{z}\rho$, $\partial_{x}T\partial_{z}T$, $\partial_{y}T\partial_{z}T$, $\partial_{z}u_{x}\partial_{x}T$, $\partial_{x}u_{z}\partial_{x}T$, $\partial_{z}u_{y}\partial_{y}T$, $\partial_{y}u_{z}\partial_{y}T$, $\partial_{z}u_{z}\partial_{y}T$, $\partial_{x}u_{x}\partial_{z}T$, $\partial_{z}u_{x}\partial_{z}T$, $\partial_{y}u_{y}\partial_{z}T$, $\partial_{z}u_{y}\partial_{z}T$, $\partial_{x}u_{z}\partial_{z}T$, $\partial_{y}u_{z}\partial_{z}T$.